OpenStudy (anonymous):

19 16/23 as a decimal ? round to the nearest tenth.

OpenStudy (anonymous):

19 16/23 = (19*23+16)/23 = 453/23 = 19.7

OpenStudy (anonymous):

- I got it, thanks love !